### v2

# Learning constitutive models and rheology from partial flow measurements

Abstract
Constitutive laws are at the core of fluid mechanics, relating the fluid stress to its deformation rate. Unlike Newtonian fluids, most industrial and biological fluids are non-Newtonian, exhibiting a nonlinear relation. Accurately characterizing this nonlinearity is essential for predicting flow behavior in real-world engineering and translational applications. Yet current methods fall short by relying on bulk rheometer data and simple fits that fail to capture behaviors relevant in complex geometries and flow conditions. Data-driven approaches can capture more complex behaviors, but lack interpretability or consistency. To close this gap, we leverage automatic differentiation to build an end-to-end framework for robust rheological learning. We develop a differentiable non-Newtonian fluid solver with a frame-invariant tensor basis neural network closure that learns stress directly from arbitrary flow measurements, such as velocimetry data. In parallel, we implement differentiable versions of major constitutive relations, enabling Bayesian model parametrization and selection from rheometer data. Our framework predicts flows in unseen geometries and ensures physical consistency and interpretability by matching neural network responses to known constitutive laws. Ultimately, this work lays the groundwork for advanced digital rheometry capable of comprehensively characterizing non-Newtonian and viscoelastic fluids under realistic in-situ or in-line operating conditions.
Keywords: rheology; automatic differentiation; inverse problems; fluid mechanics; model selection
Author Contributions: A.M.S., J.V.R., and M.P.B. designed research; A.M.S. developed the differentiable non-Newtonian fluid solver and performed the tensor basis neural network training and analysis; J.V.R. developed the differentiable model-fitting framework and performed the bulk rheometry analysis; M.G.A. and H.S.B. assisted with initial implementations and numerical validation; and A.M.S., J.V.R., and M.P.B. wrote the paper.
Modeling and understanding how fluids behave in different conditions and geometries requires an accurate mathematical description of the relationship between local stress and local deformation. Such constitutive laws serve as a pillar of fluid mechanics, materials science, and engineering, with applications ranging from drug delivery to oil extraction [24, 34, 47, 14, 49, 15]. In its most basic form, a constitutive law gives a one-dimensional functional relationship between a scalar stress and a scalar strain rate [3, 6, 48, 18]. In many cases, such simple descriptions can be extended to three dimensions to provide relationships between tensorial quantities. More complex constitutive laws can involve additional evolution equations for stress tensors [35, 36, 16]. In such cases, the constitutive equations must be solved concurrently with the Navier–Stokes equations to obtain predictive models of fluid behavior [32, 28, 10].
Figure 1: Learning rheological models from flow data. Top: A tensor basis neural network (TBNN) that maps scalar invariants of the kinematics to the stress is embedded in a differentiable flow solver. Training on any type of flow data with JAX gradients yields a TBNN that generalizes across conditions and geometries and enables flow prediction. Bottom: For interpretability and extrapolation, we also implement a differentiable ODE framework that fits the learned TBNN model’s to classical constitutive laws (e.g. Giesekus) and selects the best model and parameters using the Bayesian Information Criterion (BIC).
Different objectives motivate assigning and fitting constitutive models to fluid data [3, 11, 13]. For example, specific parameters in a simple constitutive law can aid in understanding the microstructural evolution of a fluid, such as the transition to jammed packing [27, 7]. Other constitutive laws are more phenomenological, derived mathematically to capture particular observed behaviors in distinct classes of complex fluids. Ultimately, describing a fluid in terms of any constitutive model requires inferring a set of material parameters from experimental data to enable predictive modeling in new conditions and geometries.
However, such predictions can often fail, highlighting a fundamental disconnect between how constitutive models are parameterized and how they are applied. For example, material parameters for a Boger fluid fit from simple shear rheometry often fail to predict the fluid’s strong extensional response in contractional geometries, sometimes predicting the opposite sign for the pressure drop [40, 20, 5, 19]. The causes of such failures are twofold. First, the choice of a constitutive law is itself an assumption. Parameters fit in one flow regime, such as simple shear, may mask the model’s inability to represent the physics of another, like mixed shear and extension [9, 12]. Second, standard rheological measurements, such as frequency sweeps and small-amplitude oscillatory tests, may lack sufficient kinematic detail required to constrain parameters or to distinguish between competing constitutive models [11, 33, 37, 17]. Consequently, a model may appear well-fit to rheometer data yet fail in predicting responses in other systems, not because its parameters are wrong, but because the underlying model form or the data used to fit it were not representative of the target application.
These challenges have motivated a growing body of work that moves away from classical constitutive models toward data-driven approaches for representing complex fluid behavior. One class of these methods has focused on modeling bulk rheometer data using machine learning, including Rheological Universal Differential Equations (RUDEs) [25] and Rheology Informed Neural Networks (RhINNs) [41, 8, 1, 30, 31]. In these works, the authors replace components of classical constitutive laws with neural networks trained on experimental data, which are then used to predict stresses in unseen flow conditions. Other techniques, based on sparse symbolic regression, aim to discover simple, interpretable models directly from the data [43, 42]. These data-driven methods have shown promise in fitting small data sets to produce accurate stress predictions under certain conditions. However, the machine learning models may still struggle to generalize when only limited bulk rheometry data are available for training, especially given the large number of parameters that must be inferred from sparse, bulk measurements.
Other data-driven approaches have attempted to find relevant constitutive laws from velocity measurements. One such approach that has been popular is to use Physics-Informed Neural Networks [38] to model the flow behavior[45, 39, 46, 44], but PINN-based methods learn from velocity fields with soft physics constraints and domain-specific boundary conditions, which makes them a powerful tool for model discovery but not portable across geometries. Other approaches attempt to infer constitutive laws from velocity measurements by solving adjoint-accelerated inverse problems for specific constitutive models, such as the Carreau model [23]. While powerful, these methods require formulating and solving an adjoint problem for each geometry and constitutive model considered, which can be technically demanding and computationally expensive. In addition, evaluating multiple candidate models in this framework would necessitate repeating the inversion procedure for each model separately, making broad model selection in complex geometries impractical.
In this paper, we present a novel approach that leverages automatic differentiation and differentiable simulations to efficiently learn constitutive relationships from arbitrary flow measurements and then relate these relationships back to classical constitutive laws (Fig. 1). Specifically, we develop a fully differentiable non-Newtonian fluid solver, building upon recent advances in differentiable CFD [22, 2], by embedding a tensor basis neural network (TBNN) [26] directly into the solver to learn the stress-strain relationship from experimental flow data. This approach allows us to capture the complex relationship between stress and strain in the flow without needing to perform an expensive differentiable simulation for every possible constitutive model.
We then interrogate the learned constitutive model encapsulated in the TBNN to extract interpretable physics. By fitting its stress response to established constitutive laws, we efficiently identify the classical model and parameters that best describe the learned behavior. This two-stage process transforms any flow measurement—no matter how complex the geometry—into a potential rheometer, enabling exploration of constitutive responses inaccessible in traditional experiments.
Learning rheological models from velocimetry
Figure 2: Learning a tensor basis neural network closure for stress. (a) Ground-truth steady-state $x$-velocity for pressure-driven flow through a constriction (pressure gradient $G=5$). (b) Training loss versus iteration (c) Steady-state $x$-velocity predicted by the simulation with the trained TBNN. (d) $x$-velocity at the constriction throat ($x=4$).
Learning constitutive behavior directly from complex flow fields requires a formulation that exposes how changes in material parameters propagate throughout the entire fluid domain. To this end, we developed a differentiable non-Newtonian flow solver capable of handling spatially varying viscosities (See Methods Sec. .2). Built within an automatic-differentiation framework, the solver provides exact gradients of any flow observable with respect to model parameters, enabling end-to-end optimization through the full simulation. Our immersed-boundary formulation [2] allows us to simulate arbitrary geometries and boundary conditions, enabling training directly on complex or experimentally reconstructed flow domains. This differentiable formulation thus forms the foundation for data-driven inference of constitutive laws from flow data in complex and experimentally relevant geometries.
In principle, one could recover constitutive behavior by selecting a candidate model, embedding it in the differentiable flow solver, and optimizing its parameters to match the observed flow. However, testing each model separately would require repeated expensive gradient-based simulations and is therefore computationally inefficient. Instead, we employ a tensor basis neural network (TBNN) to learn the stress–strain-rate relationship directly from data. As detailed in Methods Sec. .2, the TBNN encodes Galilean invariance by construction and, when supplied with the full invariant set and tensor bases, can represent any physically admissible stress. Thus, it serves as a geometry-agnostic constitutive representation that can be embedded in any simulation to reproduce the measured flow behavior. Conceptually, this process is illustrated in the top panel of Fig. 1, where the TBNN learns from flow data to yield a transferable constitutive model.
To demonstrate this framework, we first consider a canonical shear-thinning Carreau–Yasuda (CY) model, a standard benchmark in non-Newtonian rheology, widely used to represent the smooth shear-thinning behavior characteristic of polymer solutions and biological fluids [6, 48, 4, 29]. We use a fluid with parameters $\eta_{0}=1.0$, $\eta=0.02$, $k=5.0$, $n=0.7$, and $a=2.0$ as a baseline for validating the framework before relaxing these parameters and extending to other constitutive families. Rather than relying on simple shear, training is performed on an information-rich pressure-driven flow that spans orders of magnitude in local strain rate and thus a wide range of viscosities (Fig. S1). Specifically, we simulate flow through a constriction–expansion channel containing a semi-circular obstacle under a nondimensional pressure gradient of $G=5$ (see Methods Sec. .2 for unit definitions). The resulting steady-state velocity field, shown in Fig. 2a, exhibits strong spatial variations in shear rate that provide an ideal dataset for learning a generalized constitutive relation.
We train the TBNN on this dataset to learn the underlying stress law directly from the flow, where the neural network is formulated to predict a viscosity field as a function of local invariants (Methods Sec. .3). The model is trained using the differentiable flow solver described in Methods Sec. .2, with gradients evaluated only at steady state to minimize the loss between predicted and ground-truth velocity fields. During training, the loss decreases by more than four orders of magnitude (Fig. 2b), indicating convergence to a consistent constitutive representation. The velocity field reconstructed with the trained TBNN (Fig. 2c, S2) is visually indistinguishable from the ground truth, with the axial velocity profile at the constriction throat ($x=4$) showing near-perfect agreement. These results demonstrate that a geometry-agnostic, invariant neural representation can recover the correct constitutive mapping purely from complex flow data.
Figure 3: Flow prediction in an unseen geometry. (a) Steady-state $x$-velocity prediction for pressure-driven flow in a bidisperse porous medium with $G=7.5$. (b) Relative error compared to ground truth, binned as a function of local strain rate.
To evaluate whether the trained TBNN acts as a transferable constitutive model, we next test its ability to predict flows in conditions outside the training domain. We embed the trained model and the corresponding ground-truth CY fluid in a new geometry and increase the nondimensional pressure gradient to $G=7.5$ (Methods Sec. .2). The TBNN-predicted velocity field (Fig. 3a) closely matches the ground-truth simulation (Fig. S3). Fig. 3b shows the relative error (31) between the predicted and true velocity fields binned by local strain rate (Methods Sec. .3), demonstrating uniformly low error across all strain rates, with the smallest deviations in the range $10^{-1}<\dot{\gamma}<10^{1}$, where the training signal is strongest and most abundant.
We also verify that these results are robust to noise and degraded resolution (Methods Sec. .3), demonstrating that accuracy is maintained even with an over tenfold decrease in resolution (Fig. S4) and with correlated, heteroskedastic velocity-field noise as high as 4% of the 95th-percentile flow velocity (Figs. S5S6). Despite these degradations, the trained TBNN recovers the correct flow structure and constitutive mapping with low relative error, underscoring that a differentiable fluid solver provides strong physical regularization and enables learning even from coarse or noisy data (see Sec. Supplementary Information for detailed discussion).
Figure 4: Extracted Carreau–Yasuda parameters from a TBNN. (a) Representative oscillatory forcing: TBNN shear-stress output (points) with the best-fit Newtonian response (dashed) and Carreau–Yasuda (CY) response (solid) fit to the same trace; the CY model follows the waveform closely while the Newtonian fit misses the extrema. (b) Parity plot of the shear-thinning exponent $n$ learned from the TBNN versus ground truth across eight runs; dashed line indicates $y=x$. (c) Parity for the onset timescale $k$ on log–log axes; points fall on the identity over more than an order of magnitude. In these runs $\eta_{0}$ was fixed, and $a$ and $\eta$ are weakly constrained; numerical values for all parameters, the best-fit Newtonian viscosity, and model comparison statistics are given in Table 1. Across all runs the CY model is very strongly favored by BIC ($\Delta\mathrm{BIC}=\mathrm{BIC}_{\mathrm{N}}-\mathrm{BIC}_{\mathrm{CY}}\gg 0$).
Model interpretation and discovery from the TBNN
Beyond predictive accuracy, a key advantage of our framework is interpretability: the trained TBNN can be interrogated to reveal the constitutive behavior it has learned and to identify classical models that best describe it. To accomplish this, we employ a differentiable model-fitting method (Fig. 1, bottom panel) that probes the trained TBNN under controlled deformation histories. In this setup, we apply prescribed kinematic forcings and compute the corresponding stress response predicted by the TBNN. Because this procedure evaluates the constitutive mapping directly without solving a full flow problem, it enables rapid exploration of parameter space and extrapolation to flow conditions beyond those used in training. We then fit a library of standard rheological models to the TBNN-generated stress data and select the one that provides the best statistical agreement using the Bayesian Information Criterion (BIC).
We first apply this procedure to the TBNN trained on the constriction–expansion flow of Fig. 2. Using the digital rheometer (Methods Sec. .4), we impose sinusoidal shear-rate forcings with amplitude $f=10$ and frequency $\omega=1$ and record the corresponding stress response predicted by the TBNN. We then fit both Newtonian and Carreau–Yasuda (CY) models to the resulting stress curves (Fig. 4a). The Newtonian model fails to reproduce the nonlinear features of the response, whereas the CY model captures them accurately, yielding parameter values that closely match those used in the original simulation (first row of Table 1). Comparison of the Bayesian Information Criterion (BIC) further confirms that the CY model provides a substantially better statistical description of the learned constitutive behavior than the Newtonian model, despite having more parameters.
We then repeat this analysis for six additional TBNNs trained on fluids with varying degrees of shear thinning, using power-law exponents $n=0.6$ and $n=0.8$ to represent stronger and weaker shear-thinning behavior, respectively. For each case, we also vary the transition parameter $k$, which determines the shear rate at which thinning begins. The CY parameters recovered from the digital rheometer closely match the corresponding ground-truth values across all cases (Table 1). To visualize the consistency of parameter recovery, Fig. 4b shows the similarity between the learned and true values of $n$ and $k$, demonstrating that the TBNN accurately reproduces both the magnitude and trend of the shear-thinning response. Finally, to demonstrate that the digital rheometer is not limited to interpreting learned models but also enables direct model selection from experiments, we next apply the same tool to noisy bulk rheometer data across viscoelastic model families.
Run Source Carreau–Yasuda parameters $\eta_{N}$ $\Delta\mathrm{BIC}$
$\eta_{0}$ $\eta$ $k$ $n$ $a$
1 GT $1.000$ $0.020$ $5.000$ $0.700$ $2.000$ 0.340 915.35
Learned $0.985$ $0.021$ $4.941$ $0.701$ $1.615$
2 GT $1.000$ $0.015$ $3.000$ $0.600$ $2.000$ 0.295 814.99
Learned $0.972$ $0.017$ $3.031$ $0.615$ $2.360$
3 GT $1.000$ $0.012$ $1.000$ $0.600$ $2.000$ 0.456 434.86
Learned $0.992$ $0.013$ $0.983$ $0.639$ $1.011$
4 GT $1.000$ $0.010$ $7.000$ $0.600$ $2.000$ 0.231 221.63
Learned $1.037$ $0.010$ $6.626$ $0.602$ $2.067$
5 GT $1.000$ $0.040$ $5.000$ $0.800$ $2.000$ 0.495 526.63
Learned $1.018$ $0.040$ $4.820$ $0.792$ $2.018$
6 GT $1.000$ $0.045$ $10.000$ $0.800$ $2.000$ 0.441 670.03
Learned $1.020$ $0.045$ $9.623$ $0.793$ $2.011$
7 GT $1.000$ $0.050$ $15.000$ $0.800$ $2.000$ 0.416 555.15
Learned $1.022$ $0.050$ $14.404$ $0.795$ $2.021$
Table 1: Model extraction from TBNN. For each run, ground-truth (GT) Carreau–Yasuda (CY) parameters and CY parameters fitted to the TBNN output (Learned). $\eta_{\mathrm{N}}$ is the best-fit Newtonian viscosity to the same TBNN output. $\Delta\mathrm{BIC}\equiv\mathrm{BIC}_{\text{Newtonian}}-\mathrm{BIC}_{\text{CY}}$ (positive favors CY).
Model selection from bulk rheometry
Figure 5: Demonstration of fitting different constitutive models to the same ground-truth data. We simulate rheometer data using the same forcing functions described in the text, taking the strain rate as the controlled quantity and measuring the resulting shear stress in a Giesekus model. We fit several different constitutive models to this data and then examine how they perform when predicting the shear stress response under a new forcing that was not included in the fitting data. This new forcing and ground-truth response are shown in the top panel. In the bottom panel we show the response of the different best-fit models to this new forcing.
We now apply the same digital rheometer framework directly to noisy bulk rheometer data to identify constitutive models and recover parameters across viscoelastic families. In contrast to the TBNN setting, here the input is a time series of imposed shear rates and measured stresses, and we fit candidate constitutive laws via differentiable ODE solvers (Methods Sec. .4). This retains physical interpretability while leveraging gradient-based efficiency for statistical model comparison (BIC) under experimental noise.
While recent data-driven approaches successfully model rheometer data with neural networks, these models often rely on thousands of non-physical parameters, making them computationally expensive and difficult to interpret. Furthermore, the flow protocols probed by standard rheometric measurements are typically simple and low-dimensional, meaning that such complex models are not always warranted by the information content of the data. As an alternative, we demonstrate how combining classical, physically-grounded constitutive laws with differentiable programming provides an efficient and physically interpretable tool for fitting models to bulk shear data. Fig. 5 illustrates this workflow on synthetic rheometer data, where several candidate models are fit to a single ground-truth dataset to determine the best match. These statistical comparisons also serve as a necessary first step for assessing whether conventional constitutive models are sufficient or whether more expressive data-driven formulations are required.
To test the ability of differentiable programming to identify rheological models from data, we generated noisy synthetic rheometer data for 100 random instantiations across five common constitutive models (Newtonian, Carreau-Yasuda, Oldroyd-B, Giesekus, and Linear PTT). We then fit each dataset with all five candidate models, using a differentiable ordinary differential equation solver to perform the fitting via gradient descent (see Materials and Methods for details). We selected the best-fit model using the Bayesian Information Criterion (BIC), which penalizes models with more parameters. Our approach proved highly effective at distinguishing between the different physical models.
Figure 6: Model selection and parameter identification. (a) Confusion plot of our model-fitting approach when tested against 100 random instantiations of five common constitutive models, demonstrating that our approach achieves high fitting accuracy for all but the most complicated models. Looking at the worst fit models, we observe that when plotting the success versus failure for different parameters in (b) and (c) that the misidentification is strongly correlated with different regimes of a key parameter. We hypothesize that this is likely due to a non-optimal set of forcing functions which do not adequately sample the parameter space.
The results of this test are summarized in Fig. 6(a). Without optimizing the initial guesses or optimizer parameters for a given model, our approach is able to correctly identify the ground truth model nearly all of the time in the case of the simplest models and up to 70% of the time for the most complicated model, Linear PTT.
We can understand the failure cases in more detail by looking at where in the sampled parameter ranges our approach failed. In Figs. 6(b) and (c) we plot a key parameter from each of the two models with less than 90% identification accuracy. We see that in all three cases consistent failure, as denoted by the red symbols, is associated with strong clustering at particular values of the parameters. This indicates that our methodology cannot distinguish the underlying behavior of the model from other constitutive laws in these regimes.
While this may seem to be cause for concern, it indicates an important factor in Bayesian analysis, which is that our predictions are conditioned on the particular choice of forcing function. In this case, we generally followed the protocol for fitting rheometer-like data laid out in other data-driven approaches [25] while adding additional amplitudes to extend the range of shear stresses observed in the data. For this particular set of forcings, the underlying constitutive behavior is indistinguishable from other models, illustrating the importance of experimental design and how insufficiently informative forcings can make distinct models indistinguishable.
Table 2: Geometric median factor (estimate/true) for each parameter by model. Values <1<1 indicate median underestimation; >1>1 overestimation. For almost all models our method came within one percent of the true value. In the case of Carreau-Yasuda, our method produced reasonable parameter estimates, though some deviation may be due to forcing functions which do not adequately sample the shear-thinning regime of such models.
Newtonian Carreau-Yasuda OldroydB Giesekus LinearPTT
$\eta_{s}$ 1.000 1.000 1.000 1.000
$\eta_{p}$ 1.000 1.000 0.919
$\lambda$ 1.000 1.000 0.910
$\alpha$ 1.000
$\eta_{0}$ 1.042
$\eta$ 1.221
$a$ 0.882
$k$ 1.100
$n$ 1.006
$\zeta$ 1.001
$\epsilon$ 0.997
Once a particular model has been identified, the resulting fit can be used to estimate the material properties of the fluid. We summarize the geometric median factor (the ratio of the estimate to the true value) for each parameter by model in cases where the correct model was identified in table 2. For almost all of the models, the median factor was very nearly one, indicating low uncertainty in the fit. Most models did have a few outliers, where the predictions were off by over 100%. This generally correlated strongly with very small values of the parameters, where the learning rate of the optimizer is not tuned well to converge on these values.
While most models have parameter predictions within 10%, Carreau-Yasuda type models somewhat underperformed in terms of parameter identification. This is likely due to there being a limited range of shear rates present in our data. In particular, we can see that the zero shear rate shear-thinning exponent is fit with relatively little uncertainty but the other parameters are less accurate. Given the sinusoidal nature of our test forcings, the optimizer sees very few data points at relatively high shear rates, making it difficult to converge on specific values given the choice of optimizer settings.
For any nonlinear fitting approach, choosing the correct hyperparameters is critically important to achieve good convergence. When we repeated the Carreau-Yasuda parameter identification experiment on a new set of samples with slower hyperparameter (as described in Methods) the median error factor decreased to at most 1.003 for all parameters except for $\eta$. Our differentiable approach still systematically overestimated $\eta$ due to the lack of high shear rates sampled in the data set.
Together, these results demonstrate that differentiable model fitting can robustly identify constitutive laws and parameters when the data are informative. Just as importantly, they reveal that model identifiability is ultimately limited by the richness of the experimental protocol, underscoring the need to deliberately design forcing functions that probe the relevant rheological regimes.
Discussion and Conclusion
This work establishes a framework for learning constitutive behavior directly from flow measurements through differentiable simulations. By embedding the non-Newtonian stress within a fluid solver, we demonstrated that constitutive laws can be inferred from sparse, noisy, or indirect observations – turning rheological inference into a gradient-based optimization problem. The ability of the tensor basis neural network (TBNN) to recover the correct stress response from limited data underscores that the governing equations themselves provide a powerful inductive bias: physical structure reduces the need for large datasets and increases robustness to noise, enabling the learned data-driven constitutive model to be directly interrogated and compared against classical laws.
At the same time, our results reveal fundamental limitations of conventional bulk model fitting. Even for simple shear-thinning fluids, reliable parameter identification depends critically on the richness of the applied forcing. Classical Carreau-Yasuda fits assume shear rate sweeps spanning several orders of magnitude; when restricted to narrower protocols, parameter estimates become degenerate or misleading. This illustrates a general challenge in Bayesian model selection for rheology, where under-informative experiments, rather than model inadequacy, often dominate error. Neural network closures would require orders of magnitude more exposure to flow diversity to achieve comparable fidelity.
Because our solver is fully differentiable, it offers a path to overcome these limits through optimized experimental design. The same gradients used to fit model parameters can quantify how informative a forcing protocol is, allowing one to design inputs that maximally discriminate between competing constitutive hypotheses or minimize parameter uncertainty. In this sense, differentiable rheometry can close the loop between experiment and model, automating what has historically been an intuitive and labor-intensive process.
Beyond the velocimetry examples shown here, the framework generalizes naturally to any measurable flow quantity, such as tracer trajectories, pressure fields, flow-rate fluctuations, or combinations across conditions. This flexibility allows it to serve as a framework to turn any measurement into a rheometer. Further, our approach can enable construction of in-line or on-chip rheometers that infer constitutive behavior directly from process data, providing access to rheology in environments where sampling or laboratory rheometry would alter the material, such as emulsions, polymer melts, or multiphase suspensions. Implementing more complex physics, such as viscoelasticity and multiphase flows, is ongoing work that will expand the reach of this approach.
Looking further ahead, we envision differentiable fluid simulations combined with data-driven closures to enable model discovery when no classical constitutive law suffices: the TBNN can learn interpretable tensor bases, while symbolic regression or physical reasoning can translate them into new analytical forms. Conversely, the same differentiable infrastructure can perform inverse design, optimizing model or material parameters to achieve a target system behavior.
Together, these results mark a step toward automated rheological characterization. Differentiable simulations unify measurement, modeling, and optimization in a single framework that learns, tests, and designs models in silico. By making constitutive inference both interpretable and programmable, we envision this approach lays the foundation for data-driven yet physics-grounded exploration of complex fluids and, ultimately, for self-driving rheology experiments.
Materials and Methods
.1 Governing equations
We model incompressible, time-dependent flow with the Cauchy momentum balance and a general deviatoric stress:
$\nabla\cdot\mathbf{u}$ $=0,$ (1)
$\rho\big(\partial_{t}\mathbf{u}+(\mathbf{u}\cdot\nabla)\mathbf{u}\big)$ $=-\nabla p+\nabla\cdot\bm{\mathsf{\sigma}}+\mathbf{f},$ (2)
where $\mathbf{u}$ is velocity, $p$ pressure, $\bm{\mathsf{\sigma}}$ the deviatoric stress, and $\mathbf{f}$ an imposed body force (e.g., a uniform pressure gradient or immersed boundary force).
Constitutive laws and rheological models relate a fluid’s stress tensor $\bm{\mathsf{\sigma}}$ to the local rate-of-strain, $\bm{\mathsf{D}}=1/2(\bm{\nabla}\mathbf{u}+(\bm{\nabla}\mathbf{u})^{\mathrm{T}})$. The simplest such model is a Newtonian fluid, where the viscosity $\eta$ is a simple scalar factor relating the two,
$\bm{\mathsf{\sigma}}=2\eta\bm{\mathsf{D}}.$ (3)
[…]
$\eta(\dot{\gamma})=K\dot{\gamma}^{n-1},$ (4)
where $K$ and $n$ are the consistency and shear-thinning index, respectively. Another such model is the Carreau–Yasuda model [6, 48], given by
[…]
where $k$ is the characteristic timescale, $a$ is the transition sharpness, $n$ is the shear-thinning index, and $\eta_{0}$ and $\eta$ are the zero- and infinite-shear viscosity respectively.
When studying polymeric fluids, consisting of a long polymer dissolved in a solvent, constitutive models are written in terms of an extra stress $\bm{\mathsf{\tau}}$, such that the total fluid stress $\bm{\mathsf{\sigma}}$ is given by
$\bm{\mathsf{\sigma}}=2\eta_{s}\bm{\mathsf{D}}+\bm{\mathsf{\tau}}.$ (6)
Here, $\eta_{s}$ is called the solvent viscosity and describes the Newtonian contribution to the stress. The simplest viscoelastic model that is often used in describing fluids is the Oldroyd-B model [35], whose extra stress obeys an evolution equation
$\overset{\kern 2.04861pt\smalltriangledown}{\bm{\mathsf{\tau}}}+\lambda\bm{\mathsf{\tau}}=2\eta_{p}\bm{\mathsf{D}},$ (7)
where
$\overset{\kern 2.04861pt\smalltriangledown}{\bm{\mathsf{\tau}}}=\frac{\partial\bm{\mathsf{\tau}}}{\partial t}+\mathbf{u}\bm{\cdot}\bm{\nabla}\bm{\mathsf{\tau}}-\bm{\mathsf{\tau}}\bm{\cdot}\bm{\nabla}\mathbf{u}-\bm{\nabla}\mathbf{u}^{\mathrm{T}}\bm{\cdot}\bm{\mathsf{\tau}},$ (8)
is the upper-convected derivative, $\lambda$ is the relaxation time and $\eta_{p}$ is the polymer viscosity.
There are a number of additional terms which may be added to the Oldroyd-B model to account for additional physics. In this paper we will consider two additional models for viscoelastic polymer solutions, the Giesekus model [16] whose extra stress evolves like
$\overset{\kern 2.04861pt\smalltriangledown}{\bm{\mathsf{\tau}}}+\lambda\bm{\mathsf{\tau}}+\frac{\alpha\lambda}{\eta_{p}}\bm{\tau}\bm{\cdot}\bm{\tau}=2\eta_{p}\bm{\mathsf{D}},$ (9)
and the Linear Phan–Thien—Tanner model [36], given by
$\overset{\kern 2.04861pt\smallsquare}{\bm{\mathsf{\tau}}}+\left(1+\frac{\varepsilon\lambda}{\eta_{p}}\mathrm{tr}(\bm{\mathsf{\tau}})\right)\bm{\mathsf{\tau}}=2\eta_{p}\bm{\mathsf{D}}.$ (10)
Here,
$\overset{\kern 2.04861pt\smallsquare}{\bm{\mathsf{\tau}}}=\overset{\kern 2.04861pt\smalltriangledown}{\bm{\mathsf{\tau}}}+\zeta(\bm{\mathsf{\tau}}\bm{\cdot}\bm{\mathsf{D}}+\bm{\mathsf{D}}\bm{\cdot}\bm{\mathsf{\tau}})$ (11)
is known as the Gordon-Schowalter derivative.
.2 Differentiable non-Newtonian fluid solver
Differentiable solver and numerics
We solve the incompressible equations in a differentiable JAX-based framework that follows established components from JAX-CFD and immersed-boundary (IB) methods [22, 2]. At a high level, we use a staggered Cartesian grid, second-order central differences for diffusive terms, a conservative upwind discretization for advection, and a projection step that corrects a provisional velocity via a pressure Poisson solve to enforce $\nabla\cdot\mathbf{u}=0$.
[…]
All solver operations — state updates, pressure projection, and non-Newtonian stress evaluation — are expressed as pure JAX transformations, enabling exact reverse-mode differentiation through the full computation. Sensitivities are propagated through each iterative update without an explicit adjoint PDE derivation, yielding gradients of any scalar objective (e.g., velocity-field losses) with respect to constitutive parameters or neural network weights.
This end-to-end differentiable pipeline provides the map from constitutive parameters to flow observables and their gradients, and it is the backbone used to train the tensor basis neural network (TBNN) closure.
Solver validation
To validate the solver, we implemented classical generalized-Newtonian models, where the local viscosity is governed by the local strain rate, including the power-law and Carreau–Yasuda (CY) fluids. Both models were benchmarked against OpenFOAM simulations of steady Poiseuille flow, showing quantitative agreement in velocity and pressure profiles. All training data used in this work were thus generated by forward simulations from our differentiable solver.
Constitutive closure via a tensor basis neural network
While analytical constitutive models such as the power-law or Carreau–Yasuda form can be fit individually, doing so for every flow type or geometry rapidly becomes cumbersome and inflexible. Instead, we adopt a general tensorial representation of the stress based on a tensor basis neural network (TBNN), which expresses the stress as a sum over invariant tensorial bases weighted by scalar functions of the flow invariants [26, 25]:
[…]
where $\alpha_{i}(\mathcal{I})$ are scalar coefficient functions of an invariant set $\mathcal{I}=\{I_{1},I_{2},\dots,I_{K}\}$, and $\bm{\mathsf{B}}_{i}$ are tensor bases formed from $\nabla\mathbf{u}$ and its symmetric and antisymmetric parts. This construction guarantees frame invariance and provides a systematic, data-driven extension of classical constitutive laws.
For two-dimensional incompressible flow, the complete basis set reduces to
$\bm{\mathsf{D}}$ $=\tfrac{1}{2}\big(\nabla\mathbf{u}+\nabla\mathbf{u}\big),$ (13)
$\bm{\mathsf{W}}$ $=\tfrac{1}{2}\big(\nabla\mathbf{u}-\nabla\mathbf{u}\big),$ (14)
$I_{1}$ $=\mathrm{tr}(\bm{\mathsf{D}}^{2}),\quad I_{2}=\mathrm{tr}(\bm{\mathsf{W}}^{2}),$ (15)
$\bm{\mathsf{B}}_{1}$ $=\bm{\mathsf{D}},\quad\bm{\mathsf{B}}_{2}=\bm{\mathsf{W}}\bm{\mathsf{D}}-\bm{\mathsf{D}}\bm{\mathsf{W}}.$ (16)
Because the training data are generated from the Carreau–Yasuda model, which produces purely extensional stresses without rotation-induced components, we omit the second invariant and the antisymmetric basis. We verified this approximation by including the full set of bases and invariants in a separate training, which learned nearly zero dependence on the second invariant and antisymmetric term, confirming that their contribution is negligible. Accordingly, the simplified closure used for the final results still adheres to the TBNN framework while reducing to a generalized-Newtonian form with a data-driven viscosity:
$\bm{\mathsf{\sigma}}$ $=2\,\eta(I_{1})\,\bm{\mathsf{D}},$ (17)
We parameterize $\eta$ with a monotone, bounded head whose parameters are generated by a neural network over invariants.
[…]
$z(I_{1})$ $\equiv\log\left(\frac{\sqrt{2\,I_{1}}}{\dot{\gamma}_{\mathrm{ref}}}\right),$ (19)
[…]
where $F$ is a mixture of M=12M{=}12 logistic modes, with $\mathrm{sigm}(x)=1/(1+e^{-x})$. The mode parameters $\{\alpha_{m}(I_{1}),\mu_{m}(I_{1}),s_{m}(I_{1})\}$ are produced by a 16-unit Multi-Layer Perceptron (MLP), with $\alpha_{m}$ normalized and $s_{m}>0$, yielding a total of 646 trainable parameters. In these expressions, $\eta_{0}$ and $\eta$ denote the zero- and infinite-shear viscosities, respectively, and $\dot{\gamma}_{\mathrm{ref}}$ is a fixed reference shear-rate scale used for nondimensionalization, which we set as $\dot{\gamma}_{\mathrm{ref}}=1.0$.
This form ensures positivity of $\eta$ while allowing sufficient flexibility to capture nonlinear shear-thinning and thickening responses, providing a smooth, differentiable constitutive closure compatible with the TBNN framework.
Nondimensionalization and units
All quantities are reported in pressure-driven viscous units. Lengths are scaled by a characteristic length scale $H_{\mathrm{ref}}$, velocities by $U_{\mathrm{ref}}=G_{\mathrm{ref}}H_{\mathrm{ref}}^{2}/\eta_{0}$, times by $T_{\mathrm{ref}}=\eta_{0}/(G_{\mathrm{ref}}H_{\mathrm{ref}})$, stresses and pressures by $\tau_{\mathrm{ref}}=G_{\mathrm{ref}}H_{\mathrm{ref}}$, and shear rates by $\dot{\gamma}_{\mathrm{ref}}=G_{\mathrm{ref}}H_{\mathrm{ref}}/\eta_{0}$. Unless otherwise stated, the zero-shear viscosity is fixed at $\eta_{0}=1$ and the reference pressure gradient at $G_{\mathrm{ref}}=1$.
All training is performed in the constriction geometry, where the reference length corresponds to the gap width $H_{\mathrm{gap}}=1$, and the imposed nondimensional pressure gradient is $G^{*}=G/G_{\mathrm{ref}}=5$. As a generalization test, we additionally evaluate the trained model in a bidisperse porous-medium geometry, where $H_{\mathrm{ref}}$ is taken as the radius of the larger circular obstacle and the imposed gradient is $G^{*}=7.5$.
.3 Learning and evaluation protocol for the TBNN
Reference data and PIV emulation
Ground-truth fields are generated by forward simulations of the Carreau–Yasuda model. For Figs. 2 and 4, we train directly on the native simulation grid (no observation operator).
To mimic experimental measurements, in Figs. S4S6, we apply a PIV observation operator that first performs Hann-windowed, separable smoothing over interrogation windows and then samples at vector-center locations:
$\mathbf{y}$ $=\mathcal{S}\,\mathcal{H}\,\mathbf{u}_{\mathrm{true}}\;+\;\bm{\varepsilon},$ (21)
where $\mathcal{H}$ denotes Hann-windowed averaging and $\mathcal{S}$ denotes sampling on the coarser PIV grid.
The measurement noise $\bm{\varepsilon}$ is correlated and heteroskedastic. Correlation is introduced by a separable Gaussian smoothing on the vector grid with widths tied to the interrogation-window geometry:
$\bm{\varepsilon}$ $=\Sigma(\mathbf{x})\,\odot\,(\mathcal{G}*\bm{\xi})\;+\;\varsigma_{\mathrm{bias}}\,(\mathcal{G}*\mathbf{b}),\qquad\bm{\xi}\sim\mathcal{N}(\mathbf{0},\,\mathbf{I}),$ (22)
$\mathcal{G}$ $=\mathcal{G}(\varsigma_{x},\varsigma_{y}),\qquad\varsigma_{x}=\tfrac{1}{2}\,\chi_{\mathrm{corr}}\,\frac{W_{x}}{s_{x}},\quad\varsigma_{y}=\tfrac{1}{2}\,\chi_{\mathrm{corr}}\,\frac{W_{y}}{s_{y}},$ (23)
with $W_{x},W_{y}$ the interrogation-window sizes and $s_{x},s_{y}$ the strides (vector spacings). The nondimensional factor $\chi_{\mathrm{corr}}$ sets the correlation width of the synthetic noise relative to the vector spacing, and $\varsigma_{x},\varsigma_{y}$ are the Gaussian smoothing widths applied along each axis of the PIV vector grid.
Heteroskedasticity is modeled by a spatially varying scale field
[…]
For runs where we explicitly decreased resolution and/or added noise, we used a Hann kernel for $\mathcal{H}$, fixed $\chi_{\mathrm{corr}}=0.35$, $\beta=0.5$, and $\varsigma_{\mathrm{bias}}=0.1\,\varsigma_{\mathrm{base}}$, while varying the interrogation-window size (and thus the effective vector spacing) and the noise amplitude $\alpha$.
Training objective and schedule
We train the TBNN closure by matching the observed velocity field—either the full simulation grid or its PIV-processed counterpart—while applying mild regularization. Let $\mathbf{u}_{\mathrm{pred}}$ be the solver output and $\mathcal{O}\in\{\mathrm{Id},\,\mathcal{S}\mathcal{A}\}$ the observation operator (identity for full-grid training; downsampled window-averaging for PIV). The data-fidelity term is
[…]
where, for training directly on the simulation grid, $\mathbf{y}=\mathbf{u}_{\mathrm{true}}$.
A scale-invariant shape term emphasizes matching flow patterns independent of magnitude:
[…]
The viscosity-head prior discourages (i) excessive shear-thinning and (ii) overly bumpy profiles in log-viscosity space in order to keep the parameters such that the forward simulation is stable:
[…]
$\qquad s_{\mathrm{th}}=0.5,$ (27)
[…]
with $z=\log(\sqrt{2I_{1}}/\dot{\gamma}_{\mathrm{ref}})$ and a curvature threshold $c_{\mathrm{th}}=1.0$. In practice, the slope and curvature penalty was identically zero in nearly all runs.
The total objective is
$\mathcal{L}$ $=\lambda_{\mathrm{data}}\mathcal{L}_{\mathrm{data}}+\lambda_{\mathrm{shape}}\mathcal{L}_{\mathrm{shape}}$ (29)
$\quad+\lambda_{\mathrm{slope}}\mathcal{L}_{\mathrm{slope}}+\lambda_{\mathrm{curv}}\mathcal{L}_{\mathrm{curv}},$ (30)
where $\lambda_{\mathrm{data}}=1$, $\lambda_{\mathrm{shape}}=0.1$, $\lambda_{\mathrm{slope}}=10^{-3}$, and $\lambda_{\mathrm{curv}}=10^{-4}$.
We employ a two-stage training schedule. In the first stage, the network weights are held fixed while $\eta$ is adjusted until the sign of its gradient flips (typically within 881212 iterations). In the second stage, $\eta$ is frozen and the curvature and shape parameters of the viscosity head are trained. To avoid rapid convergence to a trivial Newtonian local minimum that lowers the loss, we fix $\eta_{0}=1$ throughout training. In practice, $\eta_{0}$ could be treated as a hyperparameter or unfrozen after convergence of the second stage. For runs 2 and 3 in Table 1, training was more stable when the mixture centers in $F(I_{1})$ were also held fixed.
Evaluation metrics
Except for the training loss, all evaluation metrics are computed against the ground-truth simulation fields, not the window-averaged or noisy observations.
In Fig. 3, we report a strain-rate–binned relative error, which quantifies the pointwise deviation normalized by the local ground-truth velocity magnitude and averaged within bins of strain rate:
[…]
In Figs. S4a and S5a, we report the domain-level relative root-mean-squared error (RRMSE), which measures the global difference between predicted and true velocity fields across all grid points:
[…]
The strain-rate–binned relative error (Fig. 3b) highlights local performance across different flow regimes, whereas the RRMSE (Figs. S4a and S5a) provides a single aggregate measure of overall predictive accuracy.
.4 Differentiable model fitting
Model simplification
The Oldroyd-B type models are all partial differential equations that depend on space and time. However, when fitting to one-dimensional shear rheometer data, we can simplify these models into ODEs following the example of [25] by assuming that since the flow should be purely azimuthal $\mathbf{u}=z\mathbf{e}_{\theta}$ while varying only in the vertical direction, the stress $\bm{\mathsf{\sigma}}$ should also have no variations in the azimuthal direction. This means that the terms $\mathbf{u}\bm{\cdot}\bm{\nabla}\bm{\mathsf{\tau}}=\mathbf{u}\bm{\cdot}\bm{\nabla}\bm{\mathsf{D}}=\mathbf{0}$ and $\bm{\mathsf{D}}$ becomes independent of any spatial coordinates, leaving the extra-stress equation as an ODE only in terms of time. This assumption inherently assumes that the Weissenberg number is small enough that out-of-plane instabilities do not develop, which would break the assumption of spatial homogeneity.
Rheometric fitting
Given that the shear rheometer forms the backbone of most laboratory experiments in complex fluids, we first explore how differentiable solvers enable fitting models to bulk rheometer data, in line with previous data-driven approaches to the same topic [25]. In particular, we assume that the Weissenberg number is small enough that the flow in the rheometer gap becomes approximately two-dimensional and spatially homogeneous. The result is that constitutive laws can be reduced to ordinary differential equations only in terms of time that are expressed in terms of the rate of strain tensor $\bm{\mathsf{D}}$ and the stress $\bm{\mathsf{\sigma}}$, with no direct appearance of the flow $\mathbf{u}$.
Under the quasi-two-dimensional flow assumption of a shear rheometer, the gradient of the flow $\bm{\nabla}\mathbf{u}=\dot{\gamma}(t)\mathbf{e}_{12}$ only has one non-zero component in the shear direction characterized by a potentially time-dependent shear rate $\dot{\gamma}$, which is spatially homogeneous across the entire rheometer. In a shear rheometer, this shear rate $\dot{\gamma}$ may be prescribed and the resulting shear stress $\sigma_{12}$ measured by averaging the torsional resistance encountered by the top plate. Alternatively, the shear stress may be prescribed and the strain measured directly. We will use the former case, though the approach would work equally well in the latter case. While the shear strain is assumed to be unidirectional, the stress $\bm{\mathsf{\sigma}}$ has six independent components which can be non-zero due to nonlinear coupling between the different stress terms in viscoelastic constitutive laws.
Under our previous assumptions, for a particular constitutive model we simulate the stress response to a particular strain rate $\dot{\gamma}$ by integrating the model forward in time to obtain all six stress terms for a given set of material parameters. While this can be done for arbitrary choices of material parameters, we are most interested in using this framework to fit models to experimental data. In this case, most experimental data takes the form of a measured shear stress, $\hat{\sigma}_{12}(t_{n})$, while the other five components of the ground truth are unknown. Here the $t_{n}$ are discrete measurement times. The standard loss function for fitting this data is the mean squared error or
[…]
where $\sigma_{12}$ is our predicted shear stress and $\theta$ the vector of material parameters that defines the model. Thus, fitting a given model to the data takes the form of a nonlinear minimization problem over $\theta$, where the minimization must be performed through the solution to the ODE which governs $\sigma_{12}$ for a given $\dot{\gamma}$.
We perform the integration using a differentiable solver Diffrax [21], which allows us to automatically take gradients of the loss function, Eq. (34), with respect to the parameters $\theta$. We can use gradient-descent algorithms to efficiently find a set of parameters $\theta$ that best describe the data by minimizing the loss.
This approach mirrors that used by neural network data-driven approaches which rely on very large (on the order over 1,000 to 10,000) numbers of parameters to fit the data, resulting in large computation times. We instead fit the data to classical constitutive models, which a handful of parameters, enabling us to efficiently fit an ensemble of potential models to the same dataset. Assuming that the errors in the fit are independent and normally distributed, the log-likelihood of the particular estimate is given by
$\hat{\mathcal{L}}_{\text{log-like}}=-\frac{N}{2}\left(\ln(2\pi\mathcal{L})+1\right),$
where $\mathcal{L}$ is the mean squared error defined in Eq. (34). The Bayesian information criterion for a particular fit to a model is then
$\text{BIC}=k\ln(N)-2\hat{\mathcal{L}}_{\text{log-like}}.$ (35)
Here, $k$ represents the number of parameters in the particular model, which is given simply as the length of the vector $\theta$. Under a Bayesian framework, the model which best describes the system given the observed data has the smallest BIC, a measure which naturally penalizes more complicated models with higher degrees of freedom.
Model fitting protocol
We apply this approach by generating 100 random models for each of Newtonian, Carreau-Yasuda, Oldroyd-B, Giesekus, and Linear PTT constitutive laws by randomly sampling material parameters for each model. The exact ranges of material parameters we used for this sample are given in Table S1. We then generate a set of ground truth data for each model by simulating synthetic shear stress data under the forcing function
$\dot{\gamma}=f\sin(\omega t)$ (36)
[…]
We reran the Carreau-Yasuda parameter identification tests on a new sample of 24 models with a learning rate of 0.001 and 50,000 epochs, which led to much better convergence onto the ground-truth parameters.
Acknowledgments
We thank Kaylie Hausknecht, Randy Ewoldt, Gareth McKinley, Dave Weitz and Kaushik Bhattacharya for important discussions. This work was supported by the Office of Naval Research (ONR N00014-23-1-2654) and the NSF AI Institute of Dynamic Systems (2112085).
Supplementary Information
S0.5 Effect of training data resolution and noise
Experimental flow measurements are typically lower in resolution and contain structured uncertainty. To assess how such factors influence learning, we synthetically degrade the training data to mimic micro-PIV conditions (Methods Sec. .3). Specifically, the velocity field is down-sampled by Hann-windowed averaging over square interrogation windows of width $W_{\mathrm{win}}$, and measurement noise $\bm{\varepsilon}$ is added according to the correlated, heteroskedastic model described in the Methods. The noise amplitude scales with the 95th-percentile velocity $U_{95}$, and higher-shear regions receive proportionally larger perturbations, reproducing the anisotropic error structure of experimental velocimetry measurements.
We first examine the effect of spatial resolution by varying $W_{\mathrm{win}}$ while keeping the noise parameters fixed. As shown in Fig. S4a, the trained TBNN maintains its predictive accuracy, quantified by the relative root-mean-squared error (RRMSEu), even when the resolution is reduced to $13\times 13$ interrogation windows. This robustness demonstrates that spatially coarse flow measurements remain highly informative, containing sufficient kinematic diversity to recover the underlying constitutive behavior far beyond what is accessible from traditional bulk rheometry. Example downsampled reconstructions are shown in Fig. S4b,c.
We next examine the effect of measurement noise using the same correlated, heteroskedastic model. Here the noise level on the horizontal axis of Fig. S5a is defined as the amplitude of the base scale $\sigma_{\mathrm{base}}=\alpha\,U_{95}$, expressed as a percentage of the 95th-percentile velocity magnitude $U_{95}$ (computed on the $29\times 29$ downsampled vector grid, where the noise is added). As the noise amplitude increases, the RRMSE between the TBNN prediction and the ground-truth field rises slightly but remains low, even for perturbations as large as 4%  $U_{95}$. Although the total loss cannot decrease as much in these cases, reflecting the irreducible mismatch introduced by measurement noise (Fig. S6a), the trained model still captures the underlying flow structure with high fidelity (Fig. S6b). This robustness underscores a key advantage of our differentiable formulation: by enforcing the governing equations as hard constraints, the model resists overfitting noisy data and instead converges to the physically consistent constitutive relation, even when the available information is degraded.
Supplementary Figures
Figure S1: Example kinematic and rheological fields used for training. (a) Local strain-rate magnitude $\log_{10}(\dot{\gamma})$ showing the broad range of kinematic conditions sampled within the constriction geometry for a shear-thinning fluid with parameters $\eta_{0}=1.0$, $\eta=0.02$, $k=5.0$, $n=0.7$, and $a=2.0$ . (b) Corresponding viscosity field $\log_{10}(\eta/\eta_{0})$ from the Carreau–Yasuda model, demonstrating a wide range of local viscosities and strong shear thinning in the throat region.
Figure S2: Velocity-field reconstruction and errors. Comparison between ground-truth and TBNN-predicted velocity components for the constriction geometry. (a,b) $y$-velocity fields showing near-perfect recovery of the cross-stream flow structure ($u_{y}$). (c,d) Absolute differences for $u_{y}$ and $u_{x}$ (the corresponding ground-truth and TBNN $u_{x}$ fields are shown in Fig. 2 of the main text), indicating very low absolute errors compared with the flow magnitude.
Figure S3: Generalization to unseen geometry. Comparison between TBNN predictions and ground-truth fields for flow through a bidisperse porous array. (a,b) Ground-truth and predicted $u_{x}$ fields. (c) Absolute error in $u_{x}$. (d–f) Analogous comparison for $u_{y}$. The learned closure transfers successfully to an out-of-training pressure drive in a geometry with different boundary conditions, preserving spatial structure and amplitude of both velocity components.
Figure S4: Robustness to coarse spatial resolution. (a) Relative root-mean-squared error (RRMSEu) of predicted velocity fields as a function of synthetic PIV resolution, with error bars showing variation across three separate runs. (b,c) Example downsampled velocity inputs used in training from the finest ($61$) and coarsest ($13$) resolutions. Despite heavy downsampling, the trained model captures the dominant flow structures and reproduces velocity magnitudes with very low global error.
Figure S5: Robustness to measurement noise. (a) RRMSEu of TBNN-predicted velocity fields versus synthetic PIV noise level, with error bars showing variation across five separate runs except the no noise case where we performed three repetitions. (b,c) Representative reconstructions for 0.5% and 4.0% noise amplitudes relative to $U_{95}$. Predictions remain stable and physically consistent even at high noise levels, confirming that the differentiable solver acts as a strong physical regularizer.
Figure S6: Training convergence and velocity-profile recovery for noisy data. (a) Evolution of the loss function during TBNN training for 4.0% noise amplitude relative to $U_{95}$ shows a smaller decrease than the noise-free example, reflecting the irreducible mismatch introduced by measurement noise. (b) Axial velocity profile at the constriction throat comparing the ground-truth field, noisy reference data, and TBNN predictions before and after training. The trained model matches the ground-truth profile with minimal residual error despite noise. Note that values within the solid obstacle (where the velocity should be zero) are excluded from training but shown here for visual continuity.
Supplementary Tables
Table S1: Parameter ranges and distributions for the random generation of ground truth models. For each parameter, we list the range of values and the distribution from which it is sampled.
Model Parameter Range Distribution
Newtonian Viscosity ($\eta$) [0.1, 10.0] Log
Carreau-Yasuda Zero-shear viscosity ($\eta_{0}$) [1.0, 100.0] Log
Infinite-shear viscosity ($\eta$) [0.01, 0.1] Log
Consistency index ($k$) [0.1, 10.0] Uniform
Power-law index ($n$) [0.2, 0.7] Uniform
Transition parameter ($a$) [0.5, 3.0] Uniform
Oldroyd-B Polymer viscosity ($\eta_{p}$) [1.0, 10.0] Uniform
Relaxation time ($\lambda$) [1.0, 10.0] Uniform
Solvent viscosity ($\eta_{s}$) [0.1, 10.0] Uniform
Giesekus Polymer viscosity ($\eta_{p}$) [0.1, 10.0] Log
Relaxation time ($\lambda$) [1.0, 10.0] Log
Solvent viscosity ($\eta_{s}$) [0.1, 10.0] Log
Mobility factor ($\alpha$) [0.01, 0.5] Uniform
Linear PTT Polymer viscosity ($\eta_{p}$) [0.1, 10.0] Log
Relaxation time ($\lambda$) [1.0, 10.0] Log
Solvent viscosity ($\eta_{s}$) [0.1, 10.0] Log
Elongational parameter ($\zeta$) [0.01, 0.2] Uniform
Mobility parameter ($\epsilon$) [0.01, 0.5] Uniform